OpenStudy (anonymous):

Solve for x if -2 = log(3x + 5)

OpenStudy (mathmale):

Question #1: What is the base of this logarithm?

OpenStudy (anonymous):

I was told if there isn't a base written out then it's 10

OpenStudy (mathmale):

that's true. So let's assume that the base is 10. Now, you have log(3x+5). How would you get rid of that log operator? Note that whatever you do to the right side of your equation, you must do to the left.

OpenStudy (anonymous):

Subtract 5?

OpenStudy (xapproachesinfinity):

to get rid of log base 10 you need to apply exponential to both sides \[10^{-2}=10^{\log (3x+5)}\]

OpenStudy (tkhunny):

Personally, I would do this, first. 3x + 5 > 0 x > -5/3 This will help you to avoid errors in the future.

OpenStudy (mathmale):

xapproachesinfinity ' s approach hinges on the fact that the log and exponential functions are inverses of one another. Note that \[10^{\log_{10}x }=x\]

OpenStudy (mathmale):

So, what's the value of \[10^{\log_{10}(3x+5) }\]

OpenStudy (mathmale):

???

OpenStudy (xapproachesinfinity):

but why would we do such thing to begin with! the reason is that \[\log_a (a^n)=n\] and also \[a^{\log_a(n)}=n\] this is from inverse relation of log and exponential

OpenStudy (anonymous):

I have no idea what the value of 10^log10(3x+5) would be

OpenStudy (anonymous):

I would I find the value?

OpenStudy (mathmale):

Two of us have explained that the log and the exponential functions are inverses of one another, and thus cancel each other out. I gave you the example \[10^{\log_{10}x }=x\]

OpenStudy (mathmale):

The expo. fn. 10^x and log to the base 10 are inverse functions and thus cancel out. What is the value of \[10^{\log_{10}(3x+5) }?\]

OpenStudy (anonymous):

So it equals 8?

OpenStudy (tkhunny):

\(\log\left((-1)^2\right) \ne 2\log(-1)\)

OpenStudy (anonymous):

So is the value 8?

OpenStudy (xapproachesinfinity):

from what we said \[10^{\log_10 (3x+5)}=3x+5\]

OpenStudy (xapproachesinfinity):

so it becomes only \[10^{-2}=3x+5 \]

OpenStudy (anonymous):

So would I find 10^-2, then solve for x?

OpenStudy (xapproachesinfinity):

\[10^{-2}=\frac{1}{100 }\] so \[\frac{1}{100}=3x+5\]

OpenStudy (xapproachesinfinity):

solve for x , this is just a the usual linear equation you are used to

OpenStudy (xapproachesinfinity):

oh actually there a problem @tkhunny has mentioned

OpenStudy (anonymous):

What problem?

OpenStudy (xapproachesinfinity):

oh nevermind... thought there is was an issue in the domain of log (3x+5)

OpenStudy (anonymous):

Okay. So in decimal form would x = -1.663333?

OpenStudy (xapproachesinfinity):

but don't worry about that, just proceed to solving the equation

OpenStudy (xapproachesinfinity):

yeah seems right to me it preferable to right your answer as compact as possible i meant use fractions. Don't use decimal unless asked to

OpenStudy (xapproachesinfinity):

write* instead of right sorry

OpenStudy (anonymous):

Yeah. I'm not good at using fractions and don't know how to convert that to a fraction.

OpenStudy (xapproachesinfinity):

the fraction is -499/300

OpenStudy (anonymous):

Alright. Thank you for all the help!

OpenStudy (xapproachesinfinity):

welcome

OpenStudy (xapproachesinfinity):

im perplexed to why @tkhunny mentioned x>-5/3

OpenStudy (xapproachesinfinity):

i mean i know for log x is defined for x>0 only for this equation there is really nothing to be cautious about

OpenStudy (tkhunny):

There is ALWAYS the DOMAIN to be cautions about.